### v1

@@ II Taxonomy of TL
TL is a method in machine learning (ML) that involves applying knowledge obtained from solving one problem to a different, yet related, problem. This part presents the mathematical framework underlying TL, beginning with the concepts of a ”domain” and a ”task.”
@@
A task $\mathcal{T}$ is distinguished by:
  • A label space $\mathcal{Y}$.
  • A goal-oriented predictive function $f(\cdot)$ that associates an input $x\in\mathcal{X}$ with an output $y\in\mathcal{Y}$. This function $f$ is generally unknown and represents the target of learning algorithms to estimate.
@@
The purpose of a TL approach is to utilize the source domain and task to improve the accuracy of the predictive function in the target domain. This involves adapting models, identifying shared feature representations, or employing other methods to make use of the similarities between the domains and tasks. Table III, summarizes the types of TL. Table IV provides a summary of various cutting-edge techniques that utilize TL types for KC, including the datasets and metrics employed.
TABLE III: Types of TL.
@@ II-A Inductive TL
II-A1 Fine-tuning
During fine-tuning, the model’s weights, initially learned from the source domain, are adjusted to better suit the target task (Fig.4(a)). This adjustment is done by continuing the training process on the target dataset, often with a smaller learning rate to make subtle changes to the weights. Fine-tuning in this context allows the model to adapt the knowledge acquired from the source domain to the specifics of the target domain and task, making it a quintessential example of inductive TL.
In [69], the authors’ research highlights the significance of fine-tuning in TL for enhancing a pre-trained VGG19 and naïve inception model’s ability to detect kidney diseases from CT images. By adapting the DL model more specifically to this task, fine-tuning significantly boosts performance, achieving an impressive accuracy in identifying various kidney diseases.
The authors in [70], focused on exploring a variety of techniques to improve model performance on enhancing the efficiency of human kidney segmentation (HKS) using U-Net architectures for computer-aided detection (CAD) and diagnosis. Among these techniques, they particularly emphasized the importance of hyperparameter fine-tuning, alongside test time data augmentations, train time data augmentations, shuffling, randomization, and normalization. This advancement underscores the critical role of meticulous data preprocessing and model optimization techniques, such as hyperparameter fine-tuning, in refining the accuracy and efficiency of DL models for precise medical image segmentation tasks.
In the study referenced by [71], the researchers introduce a smart model for classification and prediction. This model employs an enhanced deep belief network for the task of predicting diseases related to the kidneys. It incorporates fine-tuning through the back-propagation method to adjust all network parameters optimally using the up-down algorithm. This algorithm is a contrastive version of the wake-sleep algorithm, which is applied for training deep belief network with labeled datasets. The assessment of this model demonstrates its effectiveness in the early detection of chronic kidney disease and its stages, thus contributing to slowing down kidney damage progression.
In the research cited by [72], the authors introduce a DL-driven segmentation framework that emphasizes the critical role of fine-tuning. The framework operates through a two-phase process, starting with broad segmentation and moving towards more precise segmentation, leveraging advanced DL techniques. A key innovation of this framework is the adoption of a novel loss function designed to incorporate the mean and variance of pixel values specific to kidneys and renal tumors, thereby allowing for meticulous refinement. The challenge of limited data is addressed by pre-training the model on images from abdominal datasets. This initial step equips the model with the capability to discern fundamental characteristics from abdominal images, facilitated by TL. This approach underscores the importance of fine-tuning in enhancing the model’s accuracy and applicability in medical imaging.
In [53], the fine-tuning is used to involves adapting a DL model, specifically a 3D CNN called NephNet3D, to improve its ability to classify cell types in human kidney tissue using 3D nuclear staining. The fine-tuning significantly improved the model’s balanced accuracy on the new specimen’s data, demonstrating that fine-tuning NephNet3D with a small subset of labeled data from a new specimen significantly enhances its predictive performance on that specimen.
Fig.4(a) illustrates two methods of fine-tuning CNN: (i) to fine-tune the weights of the whole network and (ii) to fine-tune some certain layers. Fig.4(b) showcases a CNN instance of deep network adaptation, which modifies the distribution in the full connection layer by measuring the domain distance.
@@ II-A1 Fine-tuning
II-A2 Multi-Task Learning (MTL)
@@
II-B Transductive TL
The objective of transductive TL is to predict the labels $\{y_{j}^{T}\}_{j=1}^{n_{T}}$ for the instances in the target domain $\mathcal{D}_{T}$, leveraging the knowledge from the source domain $\mathcal{D}_{S}$.
@@ II-B Transductive TL
II-B1 Domain Adaptation
It focuses on adapting the source domain model to work effectively on the target domain by minimizing the domain shift. Techniques include feature representation alignment, domain-invariant feature learning, and discrepancy minimization between source and target domain distributions. The aim of DA is to minimize the domain shift between the source domain $\mathcal{D}_{S}$ and the target domain $\mathcal{D}_{T}$ (Fig.4(b)) [76].
@@ II-B1 Domain Adaptation
  • Feature transformation: It transforms the feature space to find a common subspace where the source and target domain data can be directly compared or combined. Methods include subspace alignment, manifold learning, and DL approaches that learn domain-agnostic feature representations [84].
@@
These strategies are not mutually exclusive and can be combined to develop more sophisticated transductive TL models that effectively utilize the knowledge from the source domain to make accurate predictions in the target domain.
In [90], the authors propose a two-stage source-free unsupervised domain-adaptation framework for cross-modality abdominal multi-organ segmentations. The proposed method only requires a well-trained source model and unlabeled target dataset and does not require access to the source data. The experimental results highlight that this approach achieves satisfactory performance for adaptation of a labeled CT dataset to an unlabeled MRI dataset on abdominal organ segmentations including the right kidney and left kidney.
In [91], the authors present a new DA approach called dimension-based disentangled dilated DA to disentangle the storage locations between the features to tackle the problem of domain shift for medical image segmentation tasks without the annotations of the target domain. They use adaptive instance normalization to encourage the content information to be stored in the spatial dimension, and the style information to be stored in the channel dimension. They validate the proposed method for cross-modality medical image segmentation tasks on public datasets such as liver, right kidney, left kidney and spleen.
II-B2 Domain Generalization
Domain generalization focuses on the challenge of developing models that can generalize well to unseen domains. In domain generalization, a model is trained on data from multiple source domains to perform effectively on data from target domains that were not encountered during training. The core objective is to learn a domain-invariant representation that captures the underlying structure of the data, enabling the model to make accurate predictions across different, unseen domains [92]. In domain generalization, the goal is to learn a model $f$ that generalizes well to any target domain $D_{T}$, given multiple source domains $\{D_{1},D_{2},\ldots,D_{K}\}$ [93]:
@@ II-B2 Domain Generalization
In [95], the authors address the challenge of single-source domain generalization in medical image segmentation, specifically noting the issue of domain shifts common in clinical image datasets. Traditional approaches, relying on global or random data augmentation, often produce samples that lack diversity and informativeness, inadequately representing the potential target domain distribution. To tackle this, the authors propose a novel data augmentation strategy for domain generalization, inspired by the inherent class-level representation invariance and style mutability of medical images. They suggest that unseen target data could be represented as a linear combination of class-specific random variables, each adhering to a location-scale distribution. This concept is operationalized through the use of constrained bézier transformation for augmenting data at both global and class-specific levels, significantly enhancing augmentation diversity. Additionally, a saliency-balancing fusion mechanism is introduced to increase the informativeness of augmented samples by utilizing gradient information for orientation and magnitude guidance. The authors theoretically demonstrate that their augmentation strategy effectively limits the generalization risk for unseen target domains. Their saliency-balancing location-scale augmentation method notably outperforms existing state-of-the-art techniques in two single-source domain generalization tasks, underscoring the potential of their approach to improve kidney and other medical image segmentation tasks by addressing the challenges posed by domain shifts.
II-B3 Self-Taught Learning
@@
II-B4 Few-Shot Learning
Few-shot learning is a subset of transductive TL that focuses on the ability of a model to learn and make accurate predictions from a very limited number of training examples. The primary challenge in few-shot learning is to design algorithms that can quickly adapt to new tasks or classes using only a few labeled samples, often referred to as ”shots.” This approach is crucial in scenarios where collecting large amounts of labeled data is impractical or impossible, such as in medical diagnosis, where rare diseases may only have a handful of known cases. Few-shot learning leverages prior knowledge, gained from related tasks or a broader dataset, to infer the new task’s structure with minimal data. Typically, In few-shot learning, the model learns from a few examples in the target domain through a meta-learning approach:
@@ II-B4 Few-Shot Learning
This involves techniques such as meta-learning, where the model learns learning strategies from multiple tasks, or embedding learning, where data is transformed into a space where tasks with little data can be more easily learned. Few-shot learning aims to mimic human learning’s efficiency and flexibility, enabling models to generalize well from limited information.
In [98], the authors address the challenge of domain adaptation in DL for medical imaging, specifically in classifying kidney stones from endoscopic images taken under various conditions. They introduce a novel approach combining few-shot learning with meta-learning. Initially, a self-supervised learning phase pre-trains the model to improve feature generalization across domains. Subsequently, meta-learning fine-tunes these features specifically for the domain of kidney stones, adapting the model to new, unseen domains effectively. This method significantly enhances the model’s generalization ability and demonstrates a promising direction for overcoming domain adaptation challenges in medical imaging applications.
The paper [99] introduces the affinity network model, an innovative approach to few-shot learning, which excels in learning from a minimal set of examples and generalizing effectively. Its core is formed by layers of k-nearest-neighbor attention pooling, an advancement of the graph attention model, versatile enough to handle various object sets, with or without an underlying graph structure. This k-nearest-neighbor attention pooling is designed to be integrated seamlessly into any neural network, akin to how convolutional layers are used. A particular subset, the feature attention layer, is adept at highlighting critical features for classification. Through testing on both artificial and real-world genomic data from the the cancer genome atlas (TCGA) database, the affinity network has demonstrated superior generalization abilities compared to conventional CNN models, especially when trained on limited KC datasets.
@@
In [101], the authors present an innovative approach for few-shot segmentation in medical imaging, specifically targeting left and right KC, to facilitate rapid adaptation of segmentation models to novel, previously unseen classes with minimal training data. They developed a segmentation framework based on episodic training within a CNN structure. Leveraging the natural spatial consistency found in medical images, a global correlation module was devised to determine the correlation between support and query images, integrating this module directly into the network. This technique is designed to enhance the network’s discriminative power, clustering features of the same class closer together while distinctly separating those of different organs. The efficacy of this method was validated using anatomical images from both CT and MRI scans, including cases related to KC.
II-C Unsupervised TL
Unsupervised TL focuses on transferring knowledge from a source domain to a target domain, where both domains lack labeled data, or the target domain lacks labeled data but the source domain may have some. The mathematical background of unsupervised TL (UTL) involves techniques and models that aim to learn useful representations or patterns that can be applied to new, unseen data. While there isn’t a single mathematical framework that encompasses all of UTL, several key concepts are commonly employed, including DA, representation learning, clustering and density estimation and autoencoders and generative models
II-C1 Representation Learning
The goal of representation learning is to find a transformation of the data that makes it easier to extract useful information for classification, regression, or other tasks. Mathematically, this involves learning a function $f:\mathcal{X}\rightarrow\mathcal{Z}$, where $\mathcal{X}$ is the input space and $\mathcal{Z}$ is a feature space that ideally captures the underlying factors of variation in the data.
In [102], the authors detail a self-supervised learning technique aimed at delineating the morphological features of molecular structures in microscopy images, critical for diagnosing disease phenotypes. They adeptly encode images of the slit diaphragm into a high-dimensional space without reliance on conventional segmentation or quantification methods. This process reveals a correlation between the self-supervised learning-generated embeddings and the slit diaphragm’s morphometric parameters, validated by established methodologies. Additionally, these embeddings mirror albuminuria levels, a key marker of kidney disease progression, and can distinguish between healthy and diseased mouse models with precision akin to traditional analyses. This self-supervised learning method facilitates insightful, label-free analysis of biomedical images, setting the stage for further refined, task-specific studies.
In [103], the authors tap into electronic health records to investigate the various patient subgroups within chronic kidney disease, applying the ConvAE open-source framework to conduct unsupervised DL on a dataset comprising 2.2 million U.S. patients from the OPTUM® EHR database. Through this methodology, they employ representation learning to create numerical representations of patients, enabling the identification of chronic kidney disease subtypes, associated comorbidities, and demographics of rare disease groups. Additionally, by modifying an existing algorithm for assessing disease severity, they pinpoint patient subgroups prone to rapid chronic kidney disease progression. This innovative combination of representation learning and data-driven analysis offers a novel perspective on chronic kidney disease diversity and identifies patients at an increased risk of progressing to end-stage kidney disease, underscoring the importance of early intervention.
II-C2 Clustering and Density Estimation
Unsupervised learning often involves clustering, where the goal is to group the data into clusters that represent underlying patterns or structures. Mathematically, this can be formulated as optimizing an objective function that measures the quality of the clustering. For example, the k-means clustering algorithm aims to minimize the within-cluster sum of squares:
$\min_{\{\mathbf{c}_{k}\}}\sum_{k=1}^{K}\sum_{\mathbf{x}_{i}\in C_{k}}\|\mathbf{x}_{i}-\mathbf{c}_{k}\|^{2}$
where $C_{k}$ is the set of points assigned to cluster $k$ and $\mathbf{c}_{k}$ is the centroid of cluster $k$.
Density estimation, such as using gaussian mixture models, involves estimating the probability distribution that generated the observed data. This can be particularly useful for capturing the complex distributions of data in both source and target domains.
In [104], the authors detail a method to assess kidney function loss after renal transplants by monitoring monthly estimated glomerular filtration rate changes over two years. The approach involves principal component analysis to categorize the estimated glomerular filtration rate data into two clusters, followed by the use of kernel density estimation to analyze the distribution of estimated glomerular filtration rate values within these clusters. This reveals distinct patterns in kidney function post-transplant, important for patient monitoring and treatment adjustments.
In [105], digital image processing techniques were applied to 3D CT scans to identify and diagnose various kidney disorders, including stones, cancer (tumors), cysts, and renal fibrosis, alongside healthy kidney tissue. Utilizing k-means clustering and co-occurrence matrices for segmentation and texture analysis, this method effectively distinguishes between healthy and diseased tissues, enabling the detection of tumors at very early stages. Additionally, density slicing transforms grayscale images into color-coded images based on texture density, aiding in the visualization of different tissue conditions. The analysis revealed that KC tissues exhibit distinct statistical properties like higher standard deviation and mean values, contrasting significantly from other conditions, thus facilitating early diagnosis through detailed tissue characterization.
This study [106], addresses chronic kidney disease, characterized by a progressive decline in kidney function, commonly caused by diabetes and high blood pressure. It aims to enhance diagnosis accuracy and efficiency by applying clustering algorithms to patient test data. The research focuses on categorizing chronic kidney disease stages by severity using algorithms such as k-means, k-medoids, and Fuzzy C Means. Findings indicate that the Fuzzy C Means algorithm outperforms others in distinguishing between different chronic kidney disease stages, thereby promising quicker and more precise diagnoses.
In this study [107], examines the link between chronic kidney disease and cardiovascular risk factors within nigerian populations, specifically addressing the notable lack of research on the clustering of these conditions in black homogeneous groups. Conducted in south-west nigeria, this research utilized a modified world health organization questionnaire for gathering a range of data and performed biochemical analysis on plasma samples from 1084 participants. Findings revealed a 14.2% prevalence of stage 3 chronic kidney disease, with an increase in hypertension and low HDL-C levels associated with greater clustering of chronic kidney disease and cardiovascular risk factors. The data indicated an inverse relationship between the estimated glomerular filtration rate and the aggregation of risk factors, highlighting the significant linkage between chronic kidney disease, proteinuria, and the clustering of cardiovascular risk factors. This study emphasizes the critical need for chronic kidney disease screening in individuals showing clusters of cardiovascular risk factors and suggests the importance of integrated management for these health issues.
II-C3 Autoencoders and Generative Models
Autoencoders learn a compressed, encoded representation of data, aiming to capture its most salient features. The autoencoder consists of an encoder function $f(\mathbf{x})$ that maps an input to a hidden representation, and a decoder function $g(\mathbf{h})$ that reconstructs the input from the hidden representation. The learning objective is to minimize the reconstruction error:
$\min_{f,g}\sum_{\mathbf{x}\in\mathcal{X}}\|\mathbf{x}-g(f(\mathbf{x}))\|^{2}$
Generative models, like variational autoencoders and generative adversarial networks, aim to model the distribution of data in such a way that new, synthetic data points can be generated. These models are powerful tools for unsupervised learning and TL, as they can capture complex data distributions.
The study [108] introduces a DL framework designed to diagnose cancer by leveraging pre-trained autoencoder technologies for effective feature extraction from complex, high-dimensional gene data, with a focus on KC. By fine-tuning the model for specific tasks, the framework uses neural network classifiers to predict clinical information accurately. Notably, the model significantly outperformed traditional techniques, achieving high micro and macro f1-score (F1) across gender, race, and sample type predictions. Specifically, the introduction of a conditional variational mutation autoencoder notably enhanced race prediction accuracy. The findings underscore the framework’s potential in improving prognosis, prevention, and early detection of KC.
TABLE IV: A summary of TL types for KC.
@@ II-C3 Autoencoders and Generative Models
Ref/Year Model-based TL TL type KC task Datasets Image features Metrics
[69] (2022) VGG19 and Naïve inception Inductive Fine-tuning Detection Private data 4000, CT ACC
[70] (2022) U-Net Inductive Fine-tuning Segmentation HuBMAP 15, CT Dice score
[72] (2024) U-Net Inductive Fine-tuning Segmentation Private data CT -
@@
[53] (2023) 3D-MS-RFCNN Inductive Fine-tuning Segmentation KiTS 1059, 512×512, CT ACC and dice score
[73] (2021) - Inductive MTL Prediction Private data CT Recall
[74] (2021) CNN Inductive MTL Classification Private data 1608, CT ACC and AUC
[75] (2018) 3D CNN Inductive MTL Segmentation Private data 203, CT Dice score
@@
[90] (2022) - Transductive DA Segmentation Private data 30, CT and 20, MRI Dice score
[91] (2023) - Transductive DA Segmentation Private data 20, CT and 20, MRI Average
[94] (2021) CNN Transductive DG Segmentation Private data 200, MRI Dice score
[95] (2023) - Transductive DG Segmentation Private data MRI Dice score
[98] (2022) DL Transductive FSL Classification Private data 600, CT ACC
[99] (2019) KNN Transductive FSL Prediction GDCD 654 samples ACC
[100] (2021) U-Net Transductive FSL Segmentation BCV 30, 3D, CT Dice score
[101] (2022) CNN Transductive FSL Segmentation Private data 3D, 20, MRI, 30, CT Dice score
[102] (2024) CNN Unsupervised RL Segmentation Private data 288, CT ACC, F1
[103] (2022) CNN Unsupervised RL Classification OPTUM®EHR US ACC
[105] (2023) K-Mean Unsupervised CDE Classification - - Mean
[106] (2017) K-Mean Unsupervised CDE Prediction UCIMLR 400 instances ACC, SEN, SPE
[108] (2022) DL Unsupervised Autoencoders Classification - - F1
III Applications of TL in KC

### v2

@@ 3 Taxonomy of TL
TL is a method in machine learning (ML) that involves applying knowledge obtained from solving one problem to a different, yet related, problem. This part presents the mathematical framework underlying TL, beginning with the concepts of a "domain" and a "task."
@@
A task $\mathcal{T}$ is distinguished by: A label space $\mathcal{Y}$; A goal-oriented predictive function $f(\cdot)$ that associates an input $x\in\mathcal{X}$ with an output $y\in\mathcal{Y}$. This function $f$ is generally unknown and represents the target of learning algorithms to estimate.
@@
The purpose of a TL approach is to utilize the source domain and task to improve the accuracy of the predictive function in the target domain. This involves adapting models, identifying shared feature representations, or employing other methods to make use of the similarities between the domains and tasks. Table 2, summarizes the types of TL. Table 3 provides a summary of various cutting-edge techniques that utilize TL types for KC, including the datasets and metrics employed.
Table 2: Types of TL.
@@ 3.1 Inductive TL
3.1.1 Fine-tuning
During fine-tuning, the model’s weights, initially learned from the source domain, are adjusted to better suit the target task (Figure 4(a)). This adjustment is done by continuing the training process on the target dataset, often with a smaller learning rate to make subtle changes to the weights. Fine-tuning in this context allows the model to adapt the knowledge acquired from the source domain to the specifics of the target domain and task, making it a quintessential example of inductive TL [50, 51, 52]. Figure 4(a) illustrates two methods of fine-tuning CNN: (i) to fine-tune the weights of the whole network and (ii) to fine-tune some certain layers. Figure 4(b) showcases a CNN instance of deep network adaptation, which modifies the distribution in the full connection layer by measuring the domain distance. For example in [53], the authors’ research highlights the significance of fine-tuning in TL for enhancing a pre-trained VGG19 and naïve inception model’s ability to detect kidney diseases from CT images. By adapting the DL model more specifically to this task, fine-tuning significantly boosts performance, achieving an impressive accuracy in identifying various kidney diseases. Similarly, the article [54] presents KidneyMultiNet, a web-based system for kidney disease detection using a hybrid ML model that combines two CNNs, DenseNet201 and Xception, and leverages TL techniques. This system, trained on a publicly available kidney CT scan dataset, achieved high performance and demonstrated the growing potential of advanced machine learning techniques in medical diagnostics. Furthermore, [55] introduces a methodology for kidney segmentation using TL with the U-Net architecture, employing a pre-trained EfficientNet as an encoder, which results in impressive segmentation precision and contributes to enhanced kidney tumor detection and treatment planning. Likewise, Kalluraya Puttur et al. [56] introduced a DL-based system for early KC detection using CT scans, building on the VGG16 architecture and enhanced with TL. This approach demonstrated the potential of DL in improving diagnostic accuracy and patient outcomes. In addition, the study [57] explores the transfer of a model trained on human MRI of ADPKD to animal models, effectively addressing challenges such as class imbalance and small datasets, highlighting the effectiveness of TL in cross-species medical imaging applications. Similarly, the study [58] introduces a TL model combining contrastive learning and ResNet-50 for renal pathology image classification, demonstrating how contrastive learning and TL can improve performance with minimal labeled data. In parallel, Sharma et al. [59] present an AI-based system using a modified InceptionV3 model for diagnosing kidney abnormalities, offering an autonomous and accurate diagnostic tool, particularly beneficial in areas with limited access to renal specialists. Additionally, Kaur et al. [60] propose a novel DL model based on EfficientNetB3 for classifying kidney diseases into four categories, demonstrating its high accuracy and practical utility in early diagnosis and treatment planning. In a similar vein, Prasher et al. [61] introduce a VGG16 TL model for diagnosing kidney stones using coronal CT images, highlighting its potential for clinical application by reducing dependence on manual diagnosis. Meanwhile, the study [62] investigates the role of activation functions in SqueezeNet’s performance for kidney disease classification, revealing that ReLU provides the best results, thereby emphasizing the influence of activation functions on medical image classification outcomes. Ghosh et al. [63] propose STREAMLINERS, a model that integrates fuzzy logic and TL-based feature extraction using DenseNet121 and ResNet101, achieving high accuracy on enhanced datasets. Similarly, Kumar et al. [64] develop a VGG16-based model that outperforms several alternatives in classifying kidney diseases, enhancing accuracy through fine-tuning and feature extraction, though it requires large labeled datasets. The study [65] proposes an image processing model for kidney stone classification using CT images, with image enhancement techniques like CLAHE and Laplacian filtering, and finds that ResNet101 performs best. This approach provides a reliable diagnostic tool for timely kidney stone detection. In the study [66], the authors introduce a smart model for kidney disease prediction, incorporating fine-tuning via a contrastive version of the wake-sleep algorithm, which effectively slows down kidney damage progression. Similarly, [67] presents a DL-driven segmentation framework that emphasizes the role of fine-tuning and novel loss functions to enhance model accuracy, particularly in kidney and renal tumor segmentation. Lastly, in [68], fine-tuning is used to adapt a 3D CNN (NephNet3D) to classify cell types in human kidney tissue, resulting in significant improvements in predictive performance.
@@ 3.1.1 Fine-tuning
3.1.2 Multi-Task learning (MTL)
@@
3.2 Transductive TL
The objective of transductive TL is to predict the labels $\{y_{j}^{T}\}_{j=1}^{n_{T}}$ for the instances in the target domain $\mathcal{D}_{T}$, leveraging the knowledge from the source domain $\mathcal{D}_{S}$, where:
@@ 3.2 Transductive TL
3.2.1 Domain adaptation
It focuses on adapting the source domain model to work effectively on the target domain by minimizing the domain shift. In other word, the aim of DA is to minimize the domain shift between the source domain $\mathcal{D}_{S}$ and the target domain $\mathcal{D}_{T}$ (Figure 4(b)) [72]. Techniques include:
@@ 3.2.1 Domain adaptation
  • Feature transformation: It transforms the feature space to find a common subspace where the source and target domain data can be directly compared or combined. Methods include:
@@
These strategies are not mutually exclusive and can be combined to develop more sophisticated transductive TL models that effectively utilize the knowledge from the source domain to make accurate predictions in the target domain. The study [85] presents reciprocal domain adaptation network for KC diagnosis, which integrates TL with mutual model adaptation and domain model learning. While mutual model adaptation extracts rich features through global and local pyramid pooling, domain model learning enhances representation learning by combining semi-supervised, domain-independent features. Consequently, reciprocal domain adaptation network effectively addresses issues like data bias, domain shifts, and privacy, providing a promising solution for early KC detection. In a similar context, Canbay et al. [86] propose a privacy-preserving TL framework, incorporating differential privacy with CNNs. Although classification accuracy decreases with increased privacy settings, the framework ensures strong performance while protecting patient data, thus offering a secure solution for early kidney disease detection. Furthermore, the study by Kwon et al. [87] focuses on improving segmentation for autosomal dominant polycystic kidney disease using U-Net and TL with MRI data. By testing different training scenarios, the study demonstrates that single-shot fast spin-echo images yield the best results, reinforcing the utility of TL in medical segmentation. Similarly, Lopez et al. [88] present a two-step TL method using ResNet50 for classifying kidney stones in endoscopic and charge-coupled device images. This approach improves accuracy and outperforms traditional models. In [89], the authors propose a two-stage source-free unsupervised domain-adaptation framework for cross-modality abdominal multi-organ segmentations. The proposed method only requires a well-trained source model and an unlabeled target dataset, and does not need access to the source data. The experimental results highlight that this approach achieves satisfactory performance in adapting a labeled CT dataset to an unlabeled MRI dataset for abdominal organ segmentations, including the right and left kidneys. Additionally, in [90], the authors present a new DA approach called dimension-based disentangled dilated DA, designed to tackle the problem of domain shift for medical image segmentation tasks without the annotations of the target domain. By using adaptive instance normalization, the method encourages the content information to be stored in the spatial dimension and the style information in the channel dimension. They validate the proposed method for cross-modality medical image segmentation tasks on public datasets, including liver, right kidney, left kidney, and spleen.
3.2.2 Domain generalization
[…]
DG focuses on the challenge of developing models that can generalize well to unseen domains. In domain generalization (DG), a model is trained on data from multiple source domains to perform effectively on data from target domains that were not encountered during training. The core objective is to learn a domain-invariant representation that captures the underlying structure of the data, enabling the model to make accurate predictions across different, unseen domains [91]. In DG, the goal is to learn a model $f$ that generalizes well to any target domain $D_{T}$, given multiple source domains $\{D_{1},D_{2},\ldots,D_{K}\}$ [92]:
@@ 3.2.2 Domain generalization
In [94], the authors address the challenge of single-source DG in medical image segmentation, specifically noting the issue of domain shifts common in clinical image datasets. Traditional approaches, relying on global or random data augmentation, often produce samples that lack diversity and informativeness, inadequately representing the potential target domain distribution. To tackle this, the authors propose a novel data augmentation strategy for DG, inspired by the inherent class-level representation invariance and style mutability of medical images. They suggest that unseen target data could be represented as a linear combination of class-specific random variables, each adhering to a location-scale distribution. This concept is operationalized through the use of constrained bézier transformation for augmenting data at both global and class-specific levels, significantly enhancing augmentation diversity. Additionally, a saliency-balancing fusion mechanism is introduced to increase the informativeness of augmented samples by utilizing gradient information for orientation and magnitude guidance. The authors theoretically demonstrate that their augmentation strategy effectively limits the generalization risk for unseen target domains. Their saliency-balancing location-scale augmentation method notably outperforms existing state-of-the-art techniques in two single-source DG tasks, underscoring the potential of their approach to improve kidney and other medical image segmentation tasks by addressing the challenges posed by domain shifts.
3.2.3 Self-Taught learning
@@
3.2.4 Few-Shot learning
[…]
FSL is a subset of transductive TL that focuses on the ability of a model to learn and make accurate predictions from a very limited number of training examples. The primary challenge in few-shot learning (FSL) is to design algorithms that can quickly adapt to new tasks or classes using only a few labeled samples, often referred to as "shots." This approach is crucial in scenarios where collecting large amounts of labeled data is impractical or impossible, such as in medical diagnosis, where rare diseases may only have a handful of known cases. FSL leverages prior knowledge, gained from related tasks or a broader dataset, to infer the new task’s structure with minimal data. Typically, In FSL, the model learns from a few examples in the target domain through a meta-learning approach:
@@ 3.2.4 Few-Shot learning
This involves techniques such as meta-learning, where the model learns learning strategies from multiple tasks, or embedding learning, where data is transformed into a space where tasks with little data can be more easily learned. FSL aims to mimic human learning’s efficiency and flexibility, enabling models to generalize well from limited information.
In [97], the authors address the challenge of DA in DL for medical imaging, specifically in classifying kidney stones from endoscopic images taken under various conditions. They introduce a novel approach combining FSL with meta-learning. Initially, a self-supervised learning phase pre-trains the model to improve feature generalization across domains. Subsequently, meta-learning fine-tunes these features specifically for the domain of kidney stones, adapting the model to new, unseen domains effectively. This method significantly enhances the model’s generalization ability and demonstrates a promising direction for overcoming DA challenges in medical imaging applications.
@@
Table 3: A summary of TL types for KC.
@@
Ref Year Model-based TL TL type KC task Datasets Image features Metrics
[53] 2022 VGG19 and Naïve inception Inductive Fine-tuning Detection Private data 4000, CT ACC
[99] 2022 U-Net Inductive Fine-tuning Segmentation HuBMAP 15, CT Dice score
[67] 2024 U-Net Inductive Fine-tuning Segmentation Private data CT -
@@
[68] 2023 3D-MS-RFCNN Inductive Fine-tuning Segmentation KiTS 1059, 512×512, CT ACC and dice score
[70] 2021 CNN Inductive MTL Classification Private data 1608, CT ACC and AUC
[71] 2018 3D CNN Inductive MTL Segmentation Private data 203, CT Dice score
[89] 2022 U-Net Transductive DA Segmentation Private data 30, CT and 20, MRI Dice score
[93] 2021 CNN Transductive DG Segmentation Private data 200, MRI Dice score
@@
[94] 2023 U-Net with EfficientNet-b2 Transductive DG Segmentation Private data MRI Dice score
[97] 2022 DL Transductive FSL Classification Private data 600, CT ACC
[100] 2019 KNN Transductive FSL Prediction GDCD 654 samples ACC
[98] 2021 U-Net Transductive FSL Segmentation BCV 30, 3D, CT Dice score
[101] 2022 CNN Transductive FSL Segmentation Private data 3D, 20, MRI, 30, CT Dice score
4 Applications of TL in KC